### v4

# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

Abstract.
Next-generation datacenters require highly efficient network load balancing to manage the growing scale of artificial intelligence (AI) training and general datacenter traffic. However, existing Ethernet-based solutions, such as Equal Cost Multi-Path (ECMP) and oblivious packet spraying (OPS), struggle to maintain high network utilization due to both increasing traffic demands and the expanding scale of datacenter topologies, which also exacerbate network failures. To address these limitations, we propose ARCANE, a lightweight decentralized per-packet adaptive load balancing algorithm designed to optimize network utilization while ensuring rapid recovery from link failures. ARCANE adapts to network conditions by caching good-performing paths. In case of a network failure, ARCANE re-routes traffic away from it in less than 100 microseconds. ARCANE is designed to be deployed with next-generation out-of-order transports, such as Ultra Ethernet, and uses less than 25 bytes of per-connection state regardless of the topology size. We extensively evaluate ARCANE in large-scale simulations and FPGA-based NICs.
submissionid: 52copyright: nonedoi: isbn: price:
1. Introduction
Modern AI training cluster networks draw on both high-performance computing (HPC) and cloud-native architectures (Hoefler et al., 2022b; Qian et al., 2024; Gangidi et al., 2024). These systems rely on Remote Direct Memory Access (RDMA) for low-latency, high-throughput communication. InfiniBand, a specialized RDMA transport, is widely used for its performance (269, 2004). More recently, operators have adopted commodity Ethernet fabrics, especially RoCEv2 (Association, 2024), which enable RDMA over standard hardware using off-the-shelf switches and NICs to reduce costs. However, as training clusters scale from around 10,000 endpoints (enough to train GPT-4 or Llama 3–scale models (Achiam et al., 2023; Dubey et al., 2024)) to over 100,000 nodes, both interconnects face fundamental limits in sustaining peak bandwidth due to:
  1. (1)
    increased traffic volume and burstiness in collective communication compared to traditional workloads (Li et al., 2020; Zhao et al., 2023),
  2. (2)
    the complexity and overhead of managing a lossless, in-order network at scale, especially under link failures and performance degradation. (Gangidi et al., 2024).
Thus, the community recognized the need for network stacks tailored to distributed training traffic while remaining compatible with commodity datacenter Ethernet infrastructure (Gangidi et al., 2024). Such proposals include SRD by Amazon (Shalev et al., 2020), Falcon by Google (goo, 2023), TTPoE by Tesla (Motors, 2024), and the upcoming Ultra Ethernet (UE) (Consortium, 2024), developed in collaboration between major tech players. Key open questions in these proposals are how to address load balancing and mitigate link failures.
Current-generation in-order Ethernet-based training systems (such as RoCEv2-based clusters) typically rely on ECMP (Hopps, 2009) or similar mechanisms for decentralized routing and load balancing. ECMP load balancing logic applies a hashing function to the 5-tuple header of each data packet to determine the next hop to take. The benefit of this scheme is that, ignoring link failures, it is unlikely to receive out-of-order packets at the destination NIC as packets belonging to the same connection will be routed through the same network path.
However, ECMP routing is fragile when different connections get hashed to the same link (Al-Fares et al., 2010; Alizadeh et al., 2014; Gangidi et al., 2024). In this scenario, flows can get hashed to the same path even when other paths are free, resulting in congestion and queue build up, which, in turn, can result in drops and go-back-N retransmissions (Hoefler et al., 2023).
Moreover, recent works have shown that link failures drastically impact both training times and economic costs. A single link failure can have 20$\approx$ 20$\times$ 20 × higher cost impact in distributed training workloads than in cloud workloads (Qian et al., 2024; Gangidi et al., 2024). This observation, along with the increasing scale these systems are growing at, highlights the need for a transport layer with a load balancing scheme that can adapt near-instantly, e.g., within a few round-trip times (RTTs), to network failures and, consequently, bandwidth asymmetries across the topology.
Several solutions have been proposed to overcome ECMP’s limitations. MPTCP (Raiciu et al., 2011), PLB (Qureshi et al., 2022), FlowBender (Kabbani et al., 2014), Flowlet Switching (Vanini et al., 2017), and Flowcell (He et al., 2015) divide flows into subflows or flowlets and then route each one individually. However, these solutions are still designed for in-order networks, making them inherently sensitive to collisions, prone to handling failures poorly, and requiring significant memory for each connection (Lu et al., 2018). Load balancers, such as Oblivious Packet Spraying (OPS) (Dixit et al., 2013) and Multi-Path RDMA (MPRDMA)  (Lu et al., 2018), that operate at a per-packet granularity, can mitigate ECMP-based collisions. However, both approaches lack mechanisms to load balance effectively in the presence of network failures. Additionally, MPRDMA is constrained by its limited support for receiving out-of-order (OOO) packets and its requirement for per-packet acknowledgments (ACKs).
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SDR (Shalev et al., 2020), UE (Lu et al., 2018), Falcon (goo, 2023) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing algorithm, Adaptive Routing with Caching and Aware Network Exploration aka ARCANE. ARCANE caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
ARCANE does not require any specific hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). ARCANE requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
We extensively evaluate ARCANE in simulation and augment the simulation findings with real hardware results in Section 4. We deploy ARCANE in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, ARCANE consistently outperforms state-of-the-art load balancing algorithms. ARCANE outperforms ECMP and OPS by up to 6×6$\times$6 × and 1.25×1.25$\times$1.25 × in symmetric networks, and by up to 4.5×4.5$\times$4.5 × and 2×2$\times$2 × in asymmetric networks, outperforming OPS by as much as 100×100$\times$100 × during short-term transient link failures.
2. ARCANE Building Blocks
In this section, we introduce the fundamental building blocks necessary to understand ARCANE logic. We first describe key congestion signals, such as ECN and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
2.1. Congestion Signals
Explicit Congestion Notification (ECN) marking:
allows switches to notify congestion by setting a bit in the traffic class field of the IP header. The receiver then sends back this marked ECN bit to the sender in its ACK packet header. The sender can choose to react to this congestion signal by adjusting its sending rate (Zhu et al., 2015; Alizadeh et al., 2010; Ye et al., 2019). Switches can employ various strategies for marking packets. For instance, in Random Early Detection (RED) (Floyd and Jacobson, 1993), switches probabilistically mark packets based on queue size, with the marking probability increasing linearly between two thresholds ($K_{min}$ and $K_{max}$).
In ARCANE, we use ECN as the congestion signal to detect path congestion due to its simplicity and widespread adoption (Hoefler et al., 2023; Nvidia, 2024). Since ECN is not marked for packets when the queue is smaller than $K_{min}$, ECN effectively filters out cases of minor queuing due to packet collisions across multiple hops, while identifying true congestion at a single bottleneck. In contrast, delay-based signals struggle to differentiate between these scenarios unless enhanced by advanced switch features, such as in-network telemetry (INT) (Wang et al., 2023).
Packet loss:
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). On current generation networks, ARCANE can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
Congestion Control (CC) algorithms:
rely on congestion signals, such as ECN marks or packet loss, to adjust a flow’s sending rate or window with the goal of maintaining high network utilization while preventing queue build up. ARCANE is designed to work well with any CC algorithm as long as they support receiving and acknowledging packets out-of-order for a given message. In particular, we show later how ARCANE works well with EQDS, a variant of DCTCP, and a proprietary internal CC algorithm (Olteanu et al., 2022; Alizadeh et al., 2010).
2.2. Load Balancing
Equal Cost Multi-Path (ECMP):
is one of the most commonly used and simple load balancing mechanisms. It works by using a hashing function to randomly choose one of the available paths for a given packet (Hopps, 2009). The hashing functions usually takes as input 5 elements (aka five tuple) from the packet header: the protocol number, source address, destination address, source port, and destination port (some variations utilize only the four tuple without the protocol number). Recent approaches have proposed incorporating additional fields, such as the Time-to-Live (TTL) or the Flow Label (in IPv6), to further refine the hash calculation (Kabbani et al., 2014; Qureshi et al., 2022).
Under normal circumstances, all packets belonging to the same flow are assigned statically to a given path since the hashing function will use the same values as input. This assignment is done statically and ignores the current network congestion and failure conditions. As a result, two or more flows might be assigned to the same path even if there are many more paths available. This will inevitably result in heavy congestion and, possibly, packet drops as a consequence. Such ECMP hash collisions are a well-documented limitation of standard ECMP (Gangidi et al., 2024; Alizadeh et al., 2014; Zhang et al., 2021).
Entropy Value (EV):
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). We leverage EVs in ARCANE to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
Entropy Values Set (EVS):
is a fixed-size set of EVs, since the number of possible values is constrained by the number of bits they can occupy in the packet header. For example, the source port field in a UDP header is assigned 16 bits, giving the EVS a size of 65536 possible values (excluding some reserved values) (Lu et al., 2018). While different numbers of bits could be allocated for the EVS, we analyze in Section 4.5.2 and Appendix B how many are required for optimal performance. It is generally advantageous for an algorithm to achieve good load balancing performance with a small EVS size since that often reduces the algorithm’s memory overhead. Note that the minimum effective EVS size also depends on the network topology: smaller topologies require fewer EVs to fully explore all available physical paths.
Oblivious Packet Spraying (OPS),:
also known as Random Packet Spraying (RPS) (Dixit et al., 2013), randomly distributes individual packets across all available paths between a sender and a receiver. This is done by selecting a random EV for every packet at the sending host or by choosing a random output port at the switches. OPS has the advantage of distributing traffic evenly across multiple paths, addressing most ECMP issues. However, it is unaware of asymmetries or failures and can still be sub-optimal even in a perfectly symmetrical network (Section 4.3.1).
3. ARCANE
ARCANE is a load balancing algorithm that relies on simple and memory-efficient endpoint logic. The key idea behind ARCANE is straightforward: when congestion is detected on a certain path, we explore alternative paths while caching and reusing paths with little to no congestion. Specifically, ARCANE uses a circular buffer of a fixed size to cache EVs of uncongested paths. By design, ARCANE can be implemented in NIC hardware or firmware with minimal memory/area footprint. ARCANE does not need any switch support besides ECMP hashing and ECN marking. Placing ARCANE at the endpoint not only simplifies deployment but also provides a broader, end-to-end view of the network, unlike fabric-side approaches which are limited to local visibility. Algorithm 1 details pseudocode for a ARCANE sender receiving an ACK and detecting a failure. Algorithm 2 describes pseudocode for a ARCANE sender sending a data packet.
3.1. Core Logic: Path Exploration and Reuse
During the first Bandwidth-Delay Product (BDP) worth of packets of a new or idle connection, a ARCANE sender explores random entropies from the EVS. This exploration is necessary because, initially, there is no fresh knowledge about the network’s state. In this warm up phase, ARCANE operates similarly to OPS.
Upon receiving a data packet, the receiver copies the EV from the received packet into the acknowledgement (ACK) packet, forwarding it back to the sender. More specifically, ACKs can use that same EV for their own header instead of using a new header field, eliminating the need for extra header space and for any changes to the packet wire format.
When an ACK arrives at the sender, if it is not ECN-marked, the EV it carries is cached in the circular buffer, and its validity bit is set to 1. Otherwise, if the ACK is ECN-marked, ARCANE does not cache the EV and discards it. When set, the validity bit indicates that an entropy has not been used after it has been added to the circular buffer. When sending a data packet out, ARCANE first checks if there are any valid EVs in its buffer. In case there is any valid EV, ARCANE reuses the oldest valid EV from the circular buffer and resets its validity bit. Otherwise, ARCANE explores a random EV from the EVS.
The circular buffer in ARCANE ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
3.2. Failure Mitigation: Freezing Mode
When a network, running without ARCANE, experiences any kind of transient (e.g., link flap) or persistent unrecoverable failure (e.g., a link or switch failure), it will take the system some time to recover from it: ranging from several milliseconds to update the ECMP routing group to several seconds if a reboot is needed, and more if a swap is needed (Anbiah and Sivalingam, 2022; Iselt et al., 2004).
If we assume that it takes $10$ ms to exclude a failed cable from a routing group, packets will still be routed to this failing group during this transient period, resulting in packet drops. Specifically, with a 4 KiB MTU and a 400 Gbps link, this could potentially result in over 120,000 packets (approximately 0.5 GB) being lost (ignoring congestion control). This becomes even more critical in the case of other failures where it takes longer to update the routing.
By default, ARCANE Core Logic would already limit utilizing failing paths since it would only recycle paths that are still active and returning ACKs without ECN. However, if the ARCANE buffer was empty, it would select a random EV that could potentially map to a failing path. To further improve the performance of ARCANE and prevent further losses, we introduce freezing mode. ARCANE enters this mode when detecting failures via indirect feedback from the network. ARCANE uses a timeout heuristic (Section 2.1) that can be enhanced with packet trimming (Appendix A) as a natural feedback from the network to detect failures along a path.
When in freezing mode, ARCANE:
  1. (1)
    avoids exploring new EVs at random since this could result in the hashing function picking a failing path,
  2. (2)
    reuses the elements that are currently in the circular buffer even if they might be invalid.
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that ARCANE will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we use random EVs to allow ARCANE to explore new paths and assess whether we detect new packets failing or not. This prevents ARCANE from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if ARCANE exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), ARCANE would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, ARCANE can be conservative and can safely enter freezing mode.
Algorithm 1 ARCANE logic upon ACK receive and failure detection.
[…]
4:
5:procedure onAck(ackPacket) $\triangleright$ Sec. 3.1
6:    if ackPacket.ecn is set then
7:        return
8:    end if
9:    if not ARCANEBuffer[head].isValid then
[…]
11:    end if
[…]
15:    if isFreezingMode and now() ¿ exitFreezingMode then $\triangleright$ Sec. 3.2
[…]
18:    end if
19:end procedure
20:
21:procedure onFailureDetection() $\triangleright$ Sec. 3.2
22:    if not isFreezingMode and exploreCounter == 0 then
[…]
25:    end if
26:end procedure
Algorithm 2 ARCANE logic on send datapath.
1$\triangleright$ Variables already listed in Algorithm 1
2:procedure getNextEV()
[…]
7:    $\triangleright$ Must be in freezing mode.
[…]
10:    end if
[…]
12:end procedure
13:
14:procedure onSend(dataPacket)
15:    if ARCANEBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) or exploreCounter then
[…]
18:    else
19:        $dataPacket.ev=\textit{getNextEV()}$()
20:    end if
21:end procedure
3.3. ARCANE Design Advantages
Simple and versatile algorithm::
The simplicity of ARCANE enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. ARCANE works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing.
Minimal NIC memory footprint::
A key advantage of ARCANE is that it does not need to track per-EV metrics and statistics. As will be discussed in Section 4.5.2, achieving good performance with OPS requires a relatively large EVS. If OPS were to maintain metrics for each EV, the memory overhead would be excessive for a hardware NIC implementation, e.g., 64 Kib to store 1 bit per entropy value for an EVS with 64K EVs. However, ARCANE only needs a fixed number of bytes in memory. More specifically, as detailed in Table 1, ARCANE requires only around $\approx 25$ bytes. Moreover, even when constrained to a small EVS, ARCANE is still able to perform well (Section 4.5.2), which can further reduce ARCANE’ memory footprint by 1 byte since Table 1 assumed 16 bits per EV.
There is a subtle observation as to why ARCANE can achieve a great performance without needing a lot of state: while the ARCANE buffer is used to cache good entropies, it is really only useful in certain scenarios like when receiving a burst of ACKs or during freezing mode. In reality, most of ARCANE’ state is on the wire, stored in the inflight data and ACK packets, which will inform ARCANE about the good paths in the network.
Component Footprint (bits)
Circular Buffer Element (× elements in buffer):
 Entropy Value  (cachedEV) 16
 Entropy Validity Bit  (isValid) 1
Global Variables:
 Head Buffer  (head) 8
 Number Valid Entropies  (numberOfValidEVs) 8
 Exit Freezing Time  (exitFreezingMode) 32
 Is Freezing Mode  (isFreezingMode) 1
 Explore Counter  (exploreCounter) 8
Total (1 elements in buffer) 74 $\approx$10 bytes
Total (8 elements in buffer) 193 $\approx$25 bytes
Table 1. Per-connection memory footprint of ARCANE.
Quick failure mitigation::
The general approach of ARCANE is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.
4. Evaluation
Our evaluation includes both large-scale simulations designed to stress test ARCANE under a variety of workloads, and experiments on real hardware at meaningful scale. The goal is to answer the following research questions:
  • Does ARCANE outperform state-of-the-art load balancers in a healthy network with a symmetric topology? (Sections 4.3.1 and 4.4.1)
  • How does ARCANE perform in the presence of topology asymmetries or when coexisting with non-ARCANE traffic? (Sections 4.3.2 and 4.4.2)
  • Can ARCANE recover quickly from failures? Is it resilient to severe or widespread failures? (Sections 4.3.3 and 4.4.3)
  • Does ARCANE maintain performance under varying network parameters and configurations? (Section 4.5)
4.1. Evaluation Setup
Baseline load balancers::
In large-scale simulations, we compare ARCANE with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
NIC congestion control::
In all simulated baseline runs, we use the same DCTCP (Alizadeh et al., 2010) variant used in MPRDMA (Lu et al., 2018). It applies per-ACK congestion window updates, allows the receiver to accept and acknowledge out-of-order packets, and reduces the congestion window by one MTU in case of packet drops. In the FPGA-based experimentation, we use a similar but proprietary CC algorithm that relies on ECN marking, congestion notification packets, and per-flow congestion window adjustments.
Network setup::
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of ARCANE), and (3) a network encountering various failures modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix D.
Simulation model::
We implement ARCANE by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
We reflect the specifications of current-generation switches in simulation parameters: a 4 KiB MTU size, a bandwidth of 400 Gbps, and a switch traversal latency of 500 ns (Broadcom, 2024; De Sensi et al., 2020). We assume uniform link lengths and latencies, with each link exhibiting a latency of 500 ns. We set the retransmission timeout (RTO) to $70\;\mu\text{s}$ which is the amount of time it takes to traverse every queue in the network if it was full plus the network-wide RTT. For each queue $K_{min}$ is set to 20% of the queue size (one BDP) and $K_{max}$ to 80% of it.
ARCANE-FPGA::
To demonstrate the effectiveness and resilience of entropy recycling in a real network environment, we also evaluate ARCANE in an end-to-end setting. We use a modified production-grade FPGA-based RDMA-capable NIC running a custom transport protocol capable of handling OOO packets by directly placing the payload in memory via RDMA. The transport implementation relies on SACKs that carry bitmaps (256-bit-wide in these experiments) relaying packet delivery updates to track completions and handle selective retransmissions. Each connection maintains an 8-entry-deep ARCANE buffer that stores UDP source ports (2B). Our experimental setup is configured to support up to 256 connections for a total memory footprint of 4KB. Given that only a single connection is active at a time, all ARCANE buffers can be hosted in a single SRAM memory and the associated logic to manage ARCANE buffers can be multiplexed across all connections, occupying negligible resources (¡0.04% of the total device logic resources).
The testbed consists of a two-tier fat-tree Ethernet network with 100G NICs and 12.8T switches. The default MTU for our FPGA NICs is 8KB and typical RTT incorporating NIC buffer delay and ACK processing through T0 and T1 are in the order of 10 and 15 us, respectively.
4.2. Workloads
We evaluate ARCANE on a mix of synthetic benchmarks, real datacenter traces, and distributed training collectives.
Synthetic benchmarks set:
consists of (1) incast, (2) permutation, and (3) tornado traffic patterns. Incast happens when multiple senders simultaneously send to one receiver. It is very common in storage workloads (Zhang et al., 2011; Chen et al., 2009) but also, with a small incast degree, in distributed training (Gangidi et al., 2024). In the permutation pattern, each node sends to a random receiver, and we ensure that each node is sending and receiving to exactly one node (Alizadeh et al., 2010). The tornado pattern is a special case of the permutation where each node sends to its ”twin” node in the other half of the tree. For example, with 128 nodes, node 0 would send to 64 and vice-versa, node 1 to 64 and so on. Tornado is an important worst case for load balancing, as each packet is required to traverse the full tree (Prisacari et al., 2013).
Datacenter traces::
We use real datacenter traces from similar previous work (Alizadeh et al., 2010; Zhang et al., 2017). We use a series of traces used for web search in production clusters. In such distribution the majority of flows are quite small (less than 100 KB) while a small number of flows are large. For each node we select randomly the receiver and run the simulation for 5 ms. More details are available in Appendix E.
AI collectives::
We show simulated results for two commonly used collectives in AI training: the Allreduce implemented via the ring and butterfly algorithm (Li et al., 2020), the Alltoall implemented using an algorithm where we limit the number of parallel connections per node ($n$ connections) (Hoefler et al., 2022a; Naumov et al., 2019).
Our baseline traffic for ARCANE-FPGA consists of 128 endpoints under two T0 switches continuously performing 4 MB ring-based AllReduce collective operations, with the logical ring laid out such that all connections traverse the T1 spine to maximize the pressure on the spine of the topology.
4.3. Simulation Results
We conduct a detailed analysis of ARCANE behavior for each network condition (see Section 4). We first examine a specific case in depth and then summarize key takeaways.
4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of ARCANE in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as ARCANE still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
Figure 1. A tornado workload with 16 MiB messages using OPS (top) and ARCANE (bottom) as load balancers.
Figure 2. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives.
Figure 3. ARCANE vs. OPS in a 32 MiB message send with an asymmetric topology.
Microscopic analysis::
We study this effect in a tornado pattern, which, theoretically, can be managed entirely by optimal load balancing. We study what happens at a TOR switch and register the link utilization of the uplinks both over the entire simulation and also at smaller time buckets. For visualization purposes we limit these runs to a 2-tier network where each switch has 8 uplinks. However, we note that this problem is present, to an even bigger degree, when using a larger number of uplinks as explained in Section 5.
In Figure 1 we visualize statistics for a single T0 switch during the workload run. In particular we show two key metrics over time: 1) on the left Y-axis we show the output port utilization at fixed time intervals of 20 us. If it goes above 400 Gbps it means that during the studied time bucket some queueing was created. If it goes below it means the output port was slightly under-utilized. 2) On the right Y-axis we show the queue size over time of the 8 output ports.
In the case of OPS, we can see that due to the random nature of it, queues are created over time and that the link utilization of each port, at small timeframes, sometimes goes significantly above and below the optimal point (15% more or less). This shows that while OPS does still a decent job at completing the workload close to the ideal completion time, it does inevitably create unpredictable queues (potentially even exceeding $K_{max}$ and causing drops) over short time periods. Such queues will cause the CC to kick in and slightly reduce the sending rate and average output port utilization.
In the same configuration with ARCANE, we notice a major difference in the bottom plot of Figure 1. In particular ARCANE converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
Looking at the port selection rate over the entire run of the workload for OPS vs. ARCANE, we observe that they are nearly equivalent. This is, again, because the problem is with the short-term collisions that OPS can experience at microscopic scale.
Macroscopic analysis::
We now focus more on the overall view comparing ARCANE with all the other state of the art algorithms in a series of benchmarks.
In Figure 2, we visualize a summary of the performance of the various algorithms by looking at the runtime of the workloads (max FCT). As expected, in the case of incast, the performance is driven almost exclusively by the CC and, hence, we do not see any major difference between all the load balancers and even ECMP performs well. However, once we move to permutation and tornado workloads, ECMP collisions start to drastically reduce the performance of ECMP. In most cases ARCANE outperforms all the other algorithms.
In the tornado case, Adaptive RoCE is able to match ARCANE since this is the ideal scenario for it: ARCANE, unlike Adaptive RoCE, still needs to guess during its initial BDP worth of packets. On the other hand, ARCANE outperforms Adaptive RoCE in the permutation pattern where taking a local optimal decision might not always lead to the best global outcome. We also see the difference between algorithms that were designed to reduce the number of out-of-order packets versus algorithms that do not have such hard constraints. Additionally, there is a distinction between algorithms that operate at packet-level granularity, such as ARCANE, OPS, BitMap, and MPRDMA, and those that operate at a coarser granularity, such as Flowlet and PLB.
For datacenter traces, we analyze the results for different load levels, ranging from 40% to 100%. Here we observe again, the clear difference between per-packet algorithms and less granular options. Even at higher load ARCANE is able to work well with a 5% advantage over OPS.
We also run distributed collectives and report their completion times. We observe how, by design, the ring AllReduce has the same performance for most of the load balancing algorithms utilized. This is because, due to its ring design, there is no opportunity for congestion to accumulate. In AllToAll, ARCANE gets an up to 20% advantage over the alternatives.
Figure 4. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
Figure 5. Synthetic traffic with ARCANE and non-ARCANE (ECMP) traffic, emulating a mixed traffic scenario.
4.3.2. Asymmetric Network Conditions
We evaluate ARCANE under different scenarios where some degree of asymmetry is created. We focus on two scenarios: 1) the network is not perfectly symmetrical because of some missing (or degraded) cables, 2) there is some background traffic in the network that is using ECMP routing, analyzing the coexistence between ARCANE and non-ARCANE traffic.
Microscopic analysis::
We visualize this problem with a simple scenario where we have a switch with $n$ input and output ports and $n$ flows active, each from a different source sending a 32 MiB message. To create an asymmetry, we reduce one of the uplinks speed to 200 Gbps while all the other links remain at 400 Gbps. In Figure 3 we visualize the output port utilization rate for OPS and ARCANE.
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, ARCANE eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for ARCANE).
Macroscopic analysis::
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where ARCANE gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load ARCANE gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll ARCANE keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
In Figure 5, we showcase one example of ARCANE sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, ARCANE: 1) shifts ARCANE traffic away from ECMP traffic in order to not slow down ARCANE traffic, 2) helps background traffic by ensuring that it will not be slowed down by ARCANE traffic. This also highlights the possibility of incrementally deploying ARCANE on ECMP-base systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 6. ARCANE vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).
Figure 7. ARCANE performance under different failure modes in a 8 MiB permutation, DC traces at 100% load and a ring AllReduce.
4.3.3. Network Failures
We focus our attention to cases where the network encounters a failure during operation. We collect data from several previous works on networking failures and also study internal logs to simulate the most commonly reported cases (Gunawi et al., 2018; Singh et al., 2020, 2021). Since the probability of a failure happening during a short simulation is low, we simulate worst case scenarios where we force individual failures to happen during the runtime of our simulations. Most particularly we focus on total or partial failures of cables and switches. We note that we limit our failures to components that would not prevent the workload from completing (single point of failure).
Microscopic analysis::
We begin with a synthetic scenario in which two TOR-switch uplink failures occur at different times: one lasting $100\,\mu\mathrm{s}$ starting at $t=100\,\mu\mathrm{s}$, and another lasting $200\,\mu\mathrm{s}$ starting at $t=350\,\mu\mathrm{s}$.
In Figure 6, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while ARCANE, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, ARCANE also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, ARCANE completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by 2.5×2.5$\times$2.5 ×. We showcase a similar analysis but for incremental and concurrent failures in Appendix D.3.
Macroscopic analysis::
We showcase three cases with a series of failure modes in Figure 7. We introduce each failure mode after a fixed interval to highlight how ARCANE dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that ARCANE provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps ARCANE stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with ARCANE increase with the amount of failures. Furthermore, random drops do not affect negatively ARCANE performance.
Figure 8. Extreme failures scenario.
To further demonstrate the resilience of ARCANE, we evaluate its performance under an extreme and unlikely scenario characterized by increasingly large and long-lasting network failures during a permutation. As shown in Figure 8, ARCANE performs close to an ideal load balancer, even with 50% of network cables failing, while PLB, the second-best alternative, lags significantly behind.
This analysis proves that ARCANE is able to guarantee good load balancing performance under short-lived failures, long-lived failures and a combination of the two even with catastrophically high failure rates.
4.4. ARCANE-FPGA Evaluation
4.4.1. Healthy Symmetric Network Conditions
The first set of results focuses on a baseline healthy symmetric network configuration. Figure 9(a) shows per-flow goodput defined as end-to-end useful bit rate observed by application, after header, overheads, retransmissions, etc. We use OPS and ARCANE across two experimental configurations: denoted as setup-1 and setup-2. In setup-1, all FPGA endpoints under the two T0s are active while in setup-2, 40 out of 64 FPGA endpoints are active.
We present results from both of these configurations as we observe small unexpected performance variations depending on which and how many switch ports are active in experiments near the peak network performance levels using all switch ports (setup-1). These variations appear to be related to internal switch microarchitectural details such as port-buffer affinity and vendor-specific scheduling policies. The setup-2 uses a subset of the switch ports and eliminates most of these vendor-specific and implementation-related behaviors. To get a better understanding of this behavior we performed a sweep where we capped the TX rate of our FPGA NICs and discovered that the slight degradation for setup-1 when using ARCANE appears to subside if the TX rate is capped at 95 Gbps.
4.4.2. Asymmetric Network Conditions
We evaluate the performance of ARCANE under asymmetric network conditions. We connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptive load balancing capabilities of ARCANE, we change the link speed of one T0-T1 link from 400 Gbps to 200 Gbps, creating asymmetry in the network. Figure 9(b) shows the per-flow goodput as observed by the application while Figure 10(a) the FCT distribution. OPS sends packets across all paths (including those crossing the 200 Gbps link) with equal probability and is ultimately capped by the slower 200 Gbps path. The ECN marking on the 200 Gbps path causes the CC algorithm to throttle all flows and eventually match the capacity of that single slower link, thus leading to underutilization of the remaining 400 Gbps links (that are running at 50% utilization).
ARCANE can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, ARCANE can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target.
((a)) Symmetric Network
((b)) Asymmetric Network
Figure 9. ARCANE-FPGA impact on goodput.
((a)) Asymmetric Network
((b)) Link Failure
Figure 10. ARCANE-FPGA impact on FCT and packet drops
4.4.3. Network Failures
We also evaluate the performance of ARCANE in the presence of network failures. To demonstrate the robustness and resilience of ARCANE in the context of link failures, Figure 10(b) shows total packet drops (average across five runs each) observed in a large-scale 128 endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), OPS continues sending packets across all paths (including those affected by the link that went down). The freezing capability of ARCANE can quickly adapt to such events (within the order of an RTO) and avoid sending packets down the affected paths as the entropy cache is replenished from packets traversing healthy paths.
4.5. ARCANE Applicability
In this section, we briefly evaluate ARCANE, in simulations, under different scenarios by changing the ACK coalescing ratio, the EVS size and, the underlying CC algorithm. Generally, we believe that ARCANE can work well under many different circumstances, topologies and workloads.
4.5.1. ACK Coalescing
We have primarily evaluated ARCANE without ACK coalescing, as this configuration allows ARCANE to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for ARCANE where the ACK only carries the entropy of the packet that triggered it.
Figure 11. Performance with different ACK coalescing ratios during a 8MiB permutation.
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, ARCANE with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), ARCANE does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
Finally, we note that the trade-off of sending more ACKs is worth the effort if the underlining hardware supports such rate and the impact on the network traffic is minimal (~1%) since ACKs are relatively small (64B) compared to the packet size usually used in modern interconnects (4KiB or 8KiB).
4.5.2. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, ARCANE, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and ARCANE when using 32, 256, and 64K EVs. ARCANE works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that ARCANE could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
4.5.3. Different CC Algorithms
ARCANE has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated ARCANE working alongside a tuned version of DCTCP. However, as we will see, ARCANE can work well even with other algorithms. Moreover, we envision also a version of ARCANE that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. ARCANE can help all of these CC algorithms when compared to OPS.
Figure 12. Performance with different EVs sizes and CCs during a 8MiB permutation.
5. Theoretical Verification
To support our experimental findings, we present a theoretical first principles analysis of OPS and demonstrate how it can lead to arbitrarily long queues. To address this issue, we develop a new recycled balls-into-bins model and prove its local convergence. This model serves as the theoretical intuition behind the ARCANE protocol (Section 3).
5.1. Recycled Balls-into-Bins Model
At maximum injection rate, OPS suffers from severe load-imbalance that eventually leads to queues of arbitrary length building up. We explain this behavior with an infinite batched balls-into-bins model (Berenbrink et al., 2018; Becchetti et al., 2019; Los and Sauerwald, 2023). In contrast, we show that recycling good paths such as in ARCANE leads to a convergent behavior and logarithmically-bounded queues.
In our switch model, each output port corresponds to a bin. At each time step, every non-empty bin removes one element. Afterward, a new set of balls (packets) arrives and is distributed among the bins. In our setting, we focus on the case where $n$ balls arrive in each time step, representing full throughput. The maximum queue length at any time step corresponds to the maximum load of any bin at that time step. Balls are removed in FIFO order from the bins.
In OPS, balls are allocated to bins uniformly at random. In what follows, we assume the EVS is sufficiently large (i.e., 16 bits), allowing us to model the assignments as uniformly random. If balls arrive at a rate of $\lambda n$ for $\lambda<1$, the process remains stable. The maximum load at any time step is $O\left(\frac{1}{1-\lambda}\log\frac{n}{1-\lambda}\right)$ with high probability (Berenbrink et al., 2018) and with probability approaching $1$, there is always a bin containing $\Omega\left(\frac{1}{1-\lambda}\log n\right)$ balls (Berenbrink et al., 2018). In the limit as $\lambda\rightarrow 1$, this implies that some bin will eventually become arbitrarily overloaded. In the context of load-balancing, this means that at the maximum injection rate, oblivious random spraying leads to unbounded queue lengths. Intuitively, this occurs because $n$ balls are introduced at each step, but fewer than $n$ may be removed, as some output ports may remain unselected.
As $n$ increases, the maximum load grows, exacerbating congestion in OPS. Figure 13 illustrates this effect for $\lambda=0.99$, where larger numbers of bins (output ports) result in faster-growing maximum queues.
Figure 13. Simulating 10000 rounds of balls-into-bins.
We propose a new model, called recycled balls-into-bins and prove it converges locally to a load-balanced state with bounded queue lengths, even at maximum injection rate.
In the recycled balls-into-bins model, we keep a set of $b\cdot n$ colors (for some constant $b$) and a threshold $\tau$. We cycle through the colors in round-robin fashion in batches of $n$ colors. In each time step, we remove a ball from each nonempty bin. If a bin has at most $\tau$ balls, the color of the removed ball remembers the bin, unless it already remembers another bin. If the bin has more than $\tau$ balls, the color forgets its bin. Then, for each color in the batch that remembers a bin, throw a ball into the remembered bin. For all colors in the batch without a remembered bin, throw a ball uniformly at random. We show that, for a single switch, recycled balls-into-bins converges, meaning all colors remember a bin and keep the same bin remembered.
Theorem 5.1.
For $n\geq 16$, $\tau\geq 4\ln n$, $b\geq 2.4\ln n$, recycled balls-into-bins converges in $O(n\log n)$ steps. Every bin has $O(\log n)$ elements throughout, with probability $1-o(1)$.
The proof in Appendix C shows that recycled balls-into-bins converges as bins fill and stabilize below the threshold.
Figure 14. Simulating 200 rounds of balls-into-bins.
To visualize these behaviors, we model the balls-into-bins problem with both OPS and with the recycled balls improvement. In Figure 14, we set $n=5$ (for visualization purposes, but this holds true for more realistic $n$ values) and showcase the queues’ evolution for OPS and the recycled balls model. The results confirm that while OPS leads to unbounded queue growth, the recycled balls model converges, keeping all queues below the threshold ($\tau$) value, consistent with our previous network simulations (e.g., Figure 1 and Figure 3).
5.2. Limitations and Alternatives
While the provided model performs well in an idealized setting with maximal injection rates, in a real network scenario, CC algorithms would reduce the sending rate if a large queue were to build up. However, such activations of the CC would result in a slowdown of the sending rate and, hence, an increase of the completion time of a workload as we can see in Section 4.3.1. Moreover, we note that while real scenarios would have more subtle interactions not modelled here (like using distinct $KMin$ and $KMax$), the key results hold and are consistent with our evaluations in Section 4. Finally, we note that here we provide a proof only about the local convergence of ARCANE and that other models such as static path assignments and round-robin across the ports would also work well and potentially provide no queueing. However, such approaches are not viable in reality for a series of reasons: they cannot react well to partial or total failures, they would struggle with multi-tier topologies and they require detailed knowledge of the underlying workloads (Gangidi et al., 2024).
6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, ARCANE aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
Generally, most load balancers work at different granularity: per-flow, sub-flow or per-packet. ECMP (Hopps, 2009) is a per-flow solution and susceptible to flow collisions and are oblivious to congested paths. Solutions like Hedera (Al-Fares et al., 2010) and MicroTE (Benson et al., 2011) require a global controller which is not desirable in production datacenters (Gangidi et al., 2024). For these reasons, many sub-flow solutions like Flowlet Switching (Vanini et al., 2017), Presto (He et al., 2015), CONGA (Alizadeh et al., 2014), PLB (Qureshi et al., 2022), FlowBender (Kabbani et al., 2014) have been proposed. However, some of these solutions are still congestion oblivious (Flowlet, Presto), react too slowly for AI/ML bursty and intense traffic (PLB, FlowBender), require specialized switches (CONGA). Moreover, most of these solutions are unable to quickly deal with blackholes and failures.
Load balancers with per-packet granularity, such as OPS (Dixit et al., 2013) or DRILL (Ghorbani et al., 2017), help drastically reduce ECMP collisions, but are still oblivious to asymmetries (for OPS) or require switch support (for DRILL). As demonstrated in Section 5 OPS can suffer even in symmetric cases. MPRDMA (Lu et al., 2018) also uses ECN for load balancing like ARCANE, however it requires probing and ACK clocking and does not offer caching of entropies. Hermes (Zhang et al., 2017) combines both ECN and delay but it works best with TCP-like protocols and has many parameters with complex tuning.
ConWeave (Song et al., 2023) is designed specifically for RDMA networks and offers a solution by masking out-of-order packets in commodity RNICs but it requires changes to TOR switches and has limited scalability.
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC (Priority Flow Control) networks while ARCANE focuses on lossy networks.
7. Conclusion
We presented ARCANE, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, ARCANE’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. ARCANE outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, ARCANE consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how ARCANE can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.
Appendix A Freezing Mode in ARCANE
Ideally, ARCANE should enter freezing mode only upon detecting a network failure. To achieve this, we employ two strategies:
  1. (1)
    Packet Trimming Support: When packet trimming is available, distinguishing between packets lost due to congestion and those lost because of network failures becomes more straightforward. We use trimming to identify and separate lost packets with greater accuracy.
  2. (2)
    Absence of Packet Trimming: In the absence of packet trimming, we analyze the maximum round-trip time (RTT) observed during a period preceding the timeout event. If the maximum RTT immediately before the timeout is high, it indicates that the packet was likely lost due to congestion. Conversely, if the maximum RTT was low, the packet was more likely lost due to a network failure.
In our paper, we focused on scenarios where packet trimming was not supported. However, ARCANE performs optimally when packet trimming is available, benefiting from both an enhanced loss detection algorithm and a more responsive CC loop.
Regardless of the employed strategy, ARCANE maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of ARCANE. As demonstrated in Section 4.5.2, ARCANE remains effective with as few as 32 EVs. For instance, we tested ARCANE with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for ARCANE with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
As an extension of ARCANE, probing can be incorporated to enhance failure detection precision. However, we chose not to integrate this into ARCANE at this stage to maintain clarity and focus on the core framework. We also do not discuss fast loss-recovery mechanisms as they are anyway orthogonal to ARCANE behaviour and could be used to further improve its performance.
Figure 15. Studying ARCANE while forcing freezing mode without failures.
Appendix B Entropy Values Set Analysis
Practical implementations make use of hashing functions to map EVs onto output ports. We show that while using a small EVS can lead to high inherent load balancing issues, using $2^{16}$ EVs is very close to uniformly random.
We first show theoretically that the EVS size is very important for OPS and later that ARCANE is much more flexible and can work well even with small EVS.
We investigate the phenomenon that a small EVS leads to poor load balancing of the output ports of a switch in the case of a fat tree topology.
We use a balls-into-bins model (Raab and Steger, 1998), where $m$ balls are thrown uniformly and idependently at random into $n$ bins. The goal is to determine the largest number of balls in any bin, referred to as the maximum load $l(m,n)$. We define the load imbalance $\lambda_{m,n}$ as $\frac{l(m,n)}{m/n}-1$, representing the extent to which the most heavily loaded bin exceeds the average.
In our setting, the output ports correspond to bins and the EVs correspond to balls. In our model, $\lambda_{m,n}$ represents the load imbalance of the EVs onto uplinks. Since EVs are chosen for packets uniformly at random, the load balancing of EVs directly affects the load balancing of packets onto output ports.
The load imbalance depends on the average number of balls per bin (i.e., $\frac{m}{n}$): if $m=n$, then the load imbalance is $\Theta\big{(}\frac{\log n}{\log\log n}\big{)}$ with high probability. However, when the ratio of balls to bins $\frac{m}{n}\gg\log n$, the load imbalance tends to zero with high probability (Raab and Steger, 1998). In conclusion, for OPS and a fixed number of flows, we expect high load imbalance with a small EVS and near-zero load imbalance as the EVS increases.
We confirm this theoretical analysis with simulations of the load imbalance. Figures 16(a) and 16(b) show the distribution of the load imbalance for $1$ and $32$ unique flows, respectively. We note that each flow is from a different sender and will hence have different header fields that will be used in the hashing function regardless of the EV value. We note that for each case, we throw for each active flow a number of balls equal to the EVS size, with each ball being a unique EV. We can see that for 32 flows, choosing less than $2^{8}$ EVs can lead to more than $10\%$ load-imbalance, whereas $2^{16}$ EVs guarantee less than a $1\%$ load imbalance.
((a)) 1 flow active
((b)) 32 flows active
Figure 16. The expected load imbalance at a switch with 32 uplinks.
Appendix C ARCANE Single Switch Convergence Proof
We prove the main theorem that demonstrates the convergence of the recycled balls-into-bins process for a single switch.
The process has two distinct phases. In the first phase, there are still empty bins. The number of balls can still grow in this ’warmup’ phase. We will bound the steps it takes to undergo the first phase and bound the maximum load. In the second phase, all bins are non-empty. Once a bin is non-empty, it remains non-empty. Moreover, once all bins are non-empty, the number of balls in the system remains constant, as for every ball removed a ball is thrown. In this second phase, we will show that the system converges to a state where all bins have at most $\tau$ balls by means of a drift theorem (Lengler, 2020). Note that once all bins have at most $\tau$ balls and every ball has a distinct color, they will keep having at most $\tau$ balls. Moreover, we will argue about the maximum load and we will show that throughout, as long as the number of colors is at least $b\cdot n\geq(2.4\ln n)n$, there is at most one ball of any given color with high probability.
Note that much of the proof is devoted to showing convergence of the second phase. If we set the threshold large enough, no bin ever overflows and essentially we only have the first phase, which converges in expected $n\ln n+n$ steps. Since queues are an expensive hardware resource, we are interested in showing good constant factor bounds for the queue lengths rather than giving a weaker asymptotic bound for $\tau$.
Proof of Theorem 5.1.
First Phase. We say a ball is fresh if it was thrown randomly. After $m=2n\ln n$ fresh balls are thrown, every bin received at least one fresh ball with probability $1-\frac{1}{n}$, by the coupon collector problem (Doerr, 2020). Either all bins are non-empty before $m$ steps and we are done, or we have thrown at least $m$ fresh balls and all bins are non-empty. Either way, a bin receives at most $O(\log n)$ fresh balls, by balls into bins with $m+n$ balls (Raab and Steger, 1998). Observe that the number of balls in a bin only increases if it receives a fresh ball. Hence, the number of received fresh balls bounds the number of balls in a bin.
Since $b>2.4\ln n$, we have at most one ball per color when the first phase ends.
Second Phase At the beginning of the second phase, we have at most $2n\ln n+n$ balls in total and all bins are nonempty. We define a potential function to measure the drift towards a completely load-balanced state. Let $X^{i}_{t}$ be the number of balls in bin $i$ at time step $t$ and let $Y^{i}_{t}=\max(0,X^{i}_{t}-\tau)$. Our potential is $Y_{t}=\sum_{i}Y^{i}_{t}$. Clearly, when the potential is 00, all bins have at most $\tau$ elements. We need to bound the drift $Y_{t+1}-Y_{t}$.
We first consider the case where each ball has a unique color. The drift depends on the number of bins with more than $\tau$ elements, which we denote by $K_{t}$. Let us consider the expected drift for one particular bin $i$. The easy cases are:
$E[Y^{i}_{t+1}-Y^{i}_{t}|X^{i}_{t}=x,K_{t}=k]=\begin{cases}\frac{k}{n}-1&\text{if }x>\tau,\\ \frac{k}{n}&\text{if }x=\tau\\ 0&\text{if }x\leq\tau-k,\\ \end{cases}$
If $x$ is in between $\tau-k$ and $\tau$, it gets interesting. The expected drift is still bounded by $\frac{n}{k}$, but this estimate is not strong enough overall. Instead, we need to show that the expected drift is very small for all such bins. To this end, we express the expected drift using a binomial random variable $B$ with parameters $\frac{1}{n}$ and $k$ as $E[\max(0,B-(\tau-x))]$. We cannot directly use linearity of expectation because of the maximum function. Instead, using total expectation leads to the following expression (for $z\geq 3$):
$E[\max(0,B-z)]$ $=\left(E\left[B\ |\ B\geq z\right]-z\right)P[B\geq z]$
$\leq\frac{1}{z-1}e^{-(z-1)},$
where we show the tail bound in Lemma C.2.
We now compute the expected drift overall by considering each type of bin separately. By summing over all bins, the total expected drift contribution of bins with less than $\tau-\ln n$ balls is therefore bounded by $\frac{1}{\ln n}$ (since $z\geq\ln n+1$). Note that these bins only contribute to the drift if $k>\ln n$. Next, consider the contribution from bins with at least $\tau-\ln n$ but at most $\tau$ balls. Since we have at most $2n\log n+n$ balls, of which at least $k(\tau+1)$ balls are in overfull bins, there can be at most $\frac{2n\log n+n-k(\tau+1)}{\tau-\ln n}\leq\frac{3}{4}n-k$ such bins. They each contribute at most $\frac{k}{n}$, so their total contribution is at most $\frac{3}{4}k-\frac{k^{2}}{n}$. Finally, each overfull bin contributes $\frac{k}{n}-1$ to the expectation, for a total contribution of $\frac{k^{2}}{n}-k$. Hence, the total expected drift is
$E[Y_{t+1}-Y_{t}|K_{t}=k]\leq\begin{cases}\frac{1}{\ln n}-\frac{k}{4}&\text{if }k>\ln n,\\ -\frac{1}{4}&\text{else.}\\ \end{cases}$
Next, we bound the expected drift when there is more than one ball per color. We have two ball of the same color only if a ball does not get removed before the next ball of the same color is thrown. This happens if more than $b$ balls arrive in the same bin in the same time step as the ball (this uses the FiFO-property of the queues). By a tail bound on the binomial distribution, this happens with probability at most $\frac{e^{5/4}}{n^{3}}$, by Lemma C.1 using $b\geq\frac{12}{5}\ln n$. By a union bound, the probability that any color in a batch has more than one ball is at most $\frac{e^{5/4}}{n^{2}}$. The drift is bounded by $n$, so the contribution to the expected drift by this case is at most $\frac{e^{5/4}}{n}$.
Note that for $n\geq 16$, $\frac{e^{5/4}}{n}+\frac{1}{\ln n}-\frac{1+\ln n}{4}\geq-\frac{1}{8}$ and $\frac{e^{5/4}}{n}-\frac{1}{4}\geq-\frac{1}{32}$. Hence, the expected drift is at most $-\frac{1}{32}$. We conclude by an additive drift theorem (Lengler, 2020; He and Yao, 2004) that the expected time until all bins are below the threshold is at most $E[32\cdot Y_{0}]=O(n\ln n)$. If all balls have unique colors at the end, the process converges then. This happens with probability $1-O(\frac{\ln n}{n})$. So with high probability, a single such $Y_{t}=0$ event suffices and the overall expectation is $O(n\ln n)$.
Observe that the number of balls in the queues is bounded by $\tau$ plus the maximum load of a batched balls-into-bins process (Berenbrink et al., 2018) with expected injection rate $\lambda=\frac{k}{n}\leq\frac{1}{2}$. Hence, the number of balls in a bin is $O(\log n)$ throughout with high probability.
We use the following tail bound:
Lemma C.1.
Let ${B}$ be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n$. Then, for any $x\geq 16$:
$P[B\geq x]\leq e^{-\frac{5}{4}(x-1)}$
Proof.
We use an additive Chernoff bound (Chung and Lu, 2006):
$P[B\geq E[B]+\delta]$ $\leq e^{-\frac{\delta^{2}}{2(E[B]+\delta/3)}}$
$\leq e^{-\frac{\delta}{\frac{2}{\delta}+\frac{2}{3}}}$
$\leq e^{-\frac{5}{4}\delta}$ using $\text{using }\delta\geq 15$
We use the following lemma to bound the expected drift of bins that are far below the threshold:
Lemma C.2.
Let ${B}$ be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n/2$. Then, for any $x\geq 7/2$:
$\left(E\left[B\ |\ B\geq x\right]-x\right)P[B\geq x]\leq\frac{1}{x-1}e^{-(x-\frac{1}{2})}$
Proof.
We use a bound from Pelekis (Pelekis, 2017) for $E\left[B\ |\ B\geq x\right]$. By their theorem on this conditional expectation:
$E\left[B\ |\ B\geq x\right]$ $\leq x+\frac{(n-x)\frac{1}{n}}{x-\frac{k}{n}+\frac{1}{n}}\leq x+\frac{1}{x-1}\enspace.$
We bound $P[B\geq x]$ similarly as in Lemma C.1:
$P[B\geq E[B]+\delta]$ $\leq e^{-\frac{\delta^{2}}{2(E[B]+\delta/3)}}$
$\leq e^{-\frac{\delta}{\frac{1}{\delta}+\frac{2}{3}}}$ using $\text{using }E[B]\leq\frac{1}{2}$
$\leq e^{-\delta}$ using $\text{using }\delta\geq 3$
Because $E[B]\leq\frac{1}{2}$, $P[B\geq x]\leq e^{-(x-\frac{1}{2})}$. ∎
Appendix D Additional Results
This section presents additional results that could not be included in the main body of the paper.
D.1. ACK Coalescing Theoretical Modeling
To assess ARCANE performance under different ACK coalescing ratios, we validate it using the theoretical model from Section 5.1.
Figure 17. Performance with different ACK coalescing ratios using the balls-into-bins models.
Figure 17 shows how the recycled balls model performs well, even with less frequent recycling. While we lose the guarantee of consistently staying below $\tau$, with 2:1 and 4:1 recycling ratios, the queues barely exceed this threshold. An 8:1 ratio still proves slightly more advantageous than OPS.
D.2. Different Tiers
We aim to verify that ARCANE performs effectively with fat-tree topologies that have three tiers. This scenario poses a slightly greater challenge for ARCANE, as a single EV must manage two hops. Nonetheless, there is no intrinsic reason why ARCANE should not perform well in such a topology.
To validate this, we execute the synthetic benchmark using the symmetric topology described in Section 4.3.1, but with three tiers instead of two. The results, shown in Figure 18, indicate that ARCANE performs comparably to the two-tier topology.
Figure 18. Studying ARCANE with a 3 tiers fat tree topology.
D.3. Incremental Failures
To further demonstrate the resilience of ARCANE under failures, we conducted an experiment where we incrementally failed all but one of a switch’s uplinks at 200 µs intervals. Figure 19 presents the permutation from the perspective of the failing switch, where three of the four uplinks were permanently disabled in a staggered manner. As expected, ARCANE enters freezing mode immediately after the first failure, ensuring that failing output ports are avoided. Notably, small utilization spikes are observed on the failing links when ARCANE exits freezing mode to verify if the failure has been resolved. Since the failures are permanent for the duration of the experiment, ARCANE promptly re-enters freezing mode after detecting unresolved issues.
In this drastic scenario, OPS performs 40× worse than ARCANE, primarily due to its inability to avoid broken links, resulting in numerous retransmissions and reduced congestion windows.
Figure 19. ARCANE vs. OPS in a 32 MiB permutation with incremental persistent failures.
Appendix E Additional Data
The datacenter traces used throughout this paper have been previously used in a number of similar works (Vanini et al., 2017; Alizadeh et al., 2010). In particular we use traces provided by Facebook and a series of traces used for web search in production clusters. The CDF distribution for such traces can be seen in Figure 20. For most of the paper we focus exclusively on the WebSearch traces.
Figure 20. CDF for different data center traces